# UNetMamba: Efficient UNet-Like Mamba for Semantic Segmentation of High-Resolution Remote Sensing Images

### v1

@@ Abstract
The semantic segmentation of high-resolution remote sensing images plays a crucial role in downstream applications such as urban planning and disaster assessment. However, existing Transformer-based methods suffer from the constraint between accuracy and efficiency. To overcome this dilemma, we propose UNetMamba, a novel Mamba-based semantic segmentation model. It incorporates a Mamba Segmentation Decoder (MSD) that can efficiently decode the complex information within high-resolution images, and a Local Supervision Module (LSM), which is train-only but can significantly enhance the perception of local contents. Extensive experiments demonstrate that UNetMamba outperforms the state-of-the-art methods with the mIoU increased by 0.87% on LoveDA and 0.36% on ISPRS Vaihingen, while achieving high efficiency through light weight, low memory footprint and low computational cost. The source code will soon be publicly available at https://github.com/EnzeZhu2001/UNetMamba.
@@ I Introduction
Driven by advancement in sensor technology, high-resolution remote sensing images have become increasingly accessible. Rich contents not only make them vital in semantic segmentation for downstream applications, but also bring great challenges in accuracy and efficiency due to high information complexity.
@@
The CNN-based method UNet[1] is a milestone in semantic segmentation, which has established the foundational architecture for subsequent works: an encoder-decoder framework with skip connections. With the success of Transformer[2] in computer vision, Wang et al.[3] introduce Transformer into the U-shape framework, significantly improving the semantic segmentation accuracy of high-resolution remote sensing images. Consequently, numerous subsequent works have focused on innovating Transformer-based methods[4, 5, 6, 7]. Despite the prosperity of Transformer in remote sensing semantic segmentation, its quadratic computational complexity and large number of parameters[8] have inevitably limited the application efficiency in high-resolution images.
@@
Recently, Mamba[9] has been proposed to address the quadratic computational complexity of Transformer[2], reducing the complexity to linear one while maintaining competitive long-distance relationship modeling capability[8]. Subsequently, Mamba-based visual foundation models[10, 11] have been designed and further applied to several visual tasks. In the field of semantic segmentation, a series of Mamba-based UNets[12, 13, 14] have achieved outstanding performance in medical image segmentation, showing the potential of visual Mamba in segmentation. For optical remote sensing image interpretation, the 2-D selective scanning mechanism of VMamba[10] has demonstrated competitive efficiency, with RSMamba[15] for classification, CDMamba[16] for change detection, and RS3Mamba[17] for semantic segmentation.
@@
To overcome the dilemma between accuracy and efficiency and then achieve efficient semantic segmentation of high-resolution remote sensing images, we propose a novel UNet-like model named UNetMamba. It consists of three main components: an encoder using ResT[18] backbone, a mamba segmentation decoder (MSD) for efficiently decoding, and a train-only local supervision module (LSM) to enhance the perception of local semantic information. The contributions of our research are summarized as follows:
@@
1) We devise MSD, a plug-and-play Mamba-based decoder for efficiently decoding semantic information from multi-scale feature maps. Specifically, VSS blocks of VMamba[10] are transferred to the decoding side, significantly reducing the parameter count while leveraging the long-distance relationship modeling capability to decode complex information.
@@
2) We also conceive a CNN-based LSM to compensate for insufficient perception of local semantic information, namely, small land-cover elements. With two different scales of convolutions and an auxiliary loss function, we enhance the perception of local semantic information through supervised training. The train-only design also saves on inference cost, further ensuring the application efficiency.
@@
3) Based on MSD and LSM, we propose UNetMamba, a UNet-like model for efficient semantic segmentation of high-resolution remote sensing images. It not only has the advantages of being lightweight and having low memory and computation cost, but also achieves state-of-the-art (SOTA) performance on two well-known high-resolution remote sensing imagery datasets.
@@ II-A Mamba Segmentation Decoder
In previous studies[12, 13, 14, 17], the benefits of employing VMamba[10] for remote sensing semantic segmentation have been established. However, models that utilize a pre-trained VMamba backbone while encoding tend to be overweight for our goal of high efficiency. Therefore, we incorporate the basic unit of VMamba named VSS block[10] into the decoding side only.
The structure of a VSS block is depicted in Fig. 1(b). The 2-D feature map $F$, which has undergone patch expansion, firstly goes through layer normalization and is then linearly embedded:
@@
Subsequently, the output splits into two information flows. One flow undergoes a 3 × 3 depth-wise convolution followed by a Silu activation function:
@@
The SS2D (2-D Selective Scan) module then scans $F^{\prime\prime}$ in four different directions, decoding the semantic information under a global receptive field and linear complexity:
@@
where the expansion and merging operation correspond to VMamba[10], while the S6 operation denotes the selective scan space state sequential model of Mamba[9]. The output of SS2D subsequently undergoes layer normalization and an element-wise multiplication with the other flow, which has also gone through the same linear embedding and Silu activation. Finally, the VSS block outputs through a linear layer with residual connection.
@@
II-B Local Supervision Module
Due to the large receptive field of VSS blocks, the MSD tends to partially overlook local semantic information while decoding . However, such details in high-resolution images are key factors to further improve the accuracy of semantic segmentation. Therefore, the LSM is proposed to enhance the perception of local details, as depicted in Fig. 1(c). Considering the relatively smaller scale of land-covers in high-resolution remote sensing scenarios, we adopt two parallel convolution branches in LSM with the kernel size set as 3 and 1, respectively. Each branch is followed by a batch normalization layer with a ReLU6 activation function:
@@ II-B Local Supervision Module
$\tilde{F^{\prime}}=\tilde{F^{\prime}}_{1}\oplus\tilde{F^{\prime}}_{3}$ (5)
@@
$\tilde{F}=\text{Upsample}\left(\text{Conv}_{1}\left(\text{Drop}\left(\tilde{F^{\prime}}\right)\right)\right)$ (6)
@@ II-C Loss Functions
To ensure the holistic performance of UNetMamba while achieving the supervisory role of LSM, loss functions utilized in training are divided into a principal one and an auxiliary one. Following prior works[3, 4, 5, 6, 7], we use a dice loss $L_{dice}$ and a cross-entropy loss $L_{ce}$ to jointly form our principal segmentation loss function:
@@
where $N$ denotes the number of samples while $K$ denotes the number of categories. $\hat{y}_{k}^{\left(n\right)}$ is the $k$-th element in one-hot encoding of the true semantic labels for sample $n$. And $y_{k}^{\left(n\right)}$ represents the confidence of sample $n$ belonging to category $k$. As to the auxiliary loss function, cross-entropy loss $L_{ce}$ is selected[7]. Therefore, the overall loss function is formulated as follows:
@@
where the weight factor $\alpha$ is set as 0.4.
@@ III-C Performance Comparison
1) Comparison on the LoveDA Dataset: As shown in Tab. I, the proposed UNetMamba achieved the best accuracy performance in quantitative comparison, with a significant improvement of 0.87% in mIoU. It is note-worthy that our UNetMamba achieved SOTA in six out of seven categories. In background and agriculture, UNetMamba led the second place by 1.61% and 2.39%, respectively. This distinct advantage indicated that, thanks to MSD, our UNetMamba gained excellent global perception ability, which made it capable of accurately segmenting large land-covers under such a high-resolution condition. The outstanding performance of UNetMamba in building, road and other small land-cover categories proved that the train-only LSM successfully enhanced the perception of local contents through supervision and generalization. The qualitative comparison results on LoveDA are illustrated in Fig. 2, which intuitively shows that the semantic segmentation mask obtained by UNetMamba is more accurate.
2) Comparison on the ISPRS Vaihingen Dataset: The proposed UNetMamba not only achieved SOTA performance on the challenging LoveDA dataset, but also performed noticeably well on the classic Vaihingen dataset. As shown in Tab. II, in terms of accuracy, our UNetMamba achieved improvements of 0.20% in mF1, 0.36% in mIoU, and 0.18% in OA, while maintaining competitive efficiency through a lightweight model (14.76M) and low cost (225.71MB, 100.52G). Fig. 3 illustrates the qualitative comparison results, with the framed areas further demonstrating the competitiveness of UNetMamba in the perception of local semantic details, which is particularly important for the semantic segmentation accuracy in high-resolution remote sensing scenes.
@@ IV Conclusion
In this paper, an efficient semantic segmentation model called UNetMamba is proposed for high-resolution remote sensing images. Considering the multi-scale land-covers and complex information in such images, our UNetMamba features a plug-and-play Mamba-based segmentation decoder MSD for efficient semantic information decoding, and a train-only local supervision module LSM that can efficiently enhance the perception of local semantic details. Extensive experiments conducted on two well-known remote sensing datasets demonstrate that UNetMamba not only outperforms other SOTA models, but also achieves lightweight and low costs. In the future, we will continue to explore the potential of novel linear modeling mechanisms in efficient interpretation of high-resolution remote sensing images.

### v2

# UNetMamba: An Efficient UNet-Like Mamba for Semantic Segmentation of High-Resolution Remote Sensing Images

@@ Abstract
Semantic segmentation of high-resolution remote sensing images is vital in downstream applications such as land-cover mapping, urban planning and disaster assessment. Existing Transformer-based methods suffer from the constraint between accuracy and efficiency, while the recently proposed Mamba is renowned for being efficient. Therefore, to overcome the dilemma, we propose UNetMamba, a UNet-like semantic segmentation model based on Mamba. It incorporates a mamba segmentation decoder (MSD) that can efficiently decode the complex information within high-resolution images, and a local supervision module (LSM), which is train-only but can significantly enhance the perception of local contents. Extensive experiments demonstrate that UNetMamba outperforms the state-of-the-art methods with mIoU increased by 0.87% on LoveDA and 0.36% on ISPRS Vaihingen, while achieving high efficiency through the lightweight design, less memory footprint and reduced computational cost. The source code is available at https://github.com/EnzeZhu2001/UNetMamba.
@@ I Introduction
Spurred by the progress in advanced imaging techniques, high-resolution remote sensing images have become increasingly accessible, which can provide richer contents than regular images. Such intricate contents render them crucial for semantic segmentation in several downstream applications, yet simultaneously present significant challenges regarding accuracy and efficiency due to the high complexity of information.
@@
As the foundation models of computer vision continuously evolve, research on semantic segmentation of high-resolution remote sensing images has flourished. The CNN-based method UNet[1] is a milestone in semantic segmentation, which has established the foundational architecture for subsequent works: an encoder-decoder framework with skip connections. With the success of Transformer[2], Wang et al.[3] introduce it into the U-shape framework, significantly improving the semantic segmentation accuracy of high-resolution remote sensing images. Consequently, numerous subsequent works have focused on innovating Transformer-based methods[4, 5, 6, 7]. Despite the prosperity of Transformer in remote sensing semantic segmentation, its quadratic computational complexity and high parameter count[8] have inevitably limited the application efficiency in high-resolution images.
@@
Recently, Mamba [9] has become a hot point. For efficiency, it successfully achieves a linear complexity, while guaranteeing accuracy through its competitive long-distance dependency modeling capability. Subsequently, Mamba-based visual foundation models[10, 11] have been well-designed. Benefiting from such models, a series of Mamba-based UNets[12, 13, 14] have achieved impressive accuracy in medical imagery segmentation, showing the potential of visual Mamba in segmentation. For remote sensing image interpretation, the 2-D selective scanning of VMamba[10] has also demonstrated promising efficiency, with RSMamba[15] for classification, CDMamba[16] for change detection, and RS3Mamba[17] for semantic segmentation. However, these pioneering works have primarily validated the effectiveness of Mamba in their respective tasks, without fully leveraging its dual advantages in accuracy and efficiency through targeted design.
@@
To overcome the dilemma between accuracy and efficiency and then achieve efficient semantic segmentation of high-resolution remote sensing images, we propose a Mamba-based UNet-like model named UNetMamba. It consists of three main components: an encoder using ResT[18] backbone, a mamba segmentation decoder (MSD) for efficiently decoding, and a neatly-designed local supervision module (LSM) to enhance the perception of local semantic information. The contributions of our research are summarized as follows:
@@
1) We devise MSD, a plug-and-play Mamba-based decoder for efficiently decoding semantic information from multi-scale feature maps. Specifically, visual state space (VSS) blocks of VMamba[10] are expressly transferred to the decoding side, significantly reducing the parameter count while leveraging the long-distance modeling capability to accurately decode complex information in a global receptive field.
@@
2) We also conceive a CNN-based LSM to address the lack of local information perception in MSD. With two different scales of convolutions and an auxiliary loss function, we enhance the perception of local semantic information through supervised training. The train-only design also saves on inference costs, further ensuring the application efficiency.
@@
3) Based on MSD and LSM, we propose UNetMamba, a UNet-like model for efficient semantic segmentation of high-resolution remote sensing images. It has the advantages of being lightweight and requiring low costs, while also achieving state-of-the-art (SOTA) performance on two well-known high-resolution remote sensing imagery datasets.
@@ II-A Mamba Segmentation Decoder
In previous studies[12, 13, 14, 17], the benefits of employing VMamba[10] for remote sensing semantic segmentation have been established. However, models that utilize a pre-trained VMamba backbone while encoding tend to be overweight for our goal of high efficiency. Therefore, we incorporate the basic unit of VMamba, namely, VSS block[10], into the decoding side only.
The structure of a VSS block is depicted in Fig. 1(b). The 2-D feature map $F$, which has undergone patch expansion, firstly goes through layer normalization and then splits into two information flows. The main flow is further linearly embedded as $F^{\prime}$:
@@
Subsequently, $F^{\prime}$ undergoes a 3 × 3 depth-wise convolution and is then SiLU-activated into $F^{\prime\prime}$:
@@
The 2-D Selective Scan (SS2D) module thereafter scans $F^{\prime\prime}$ in four different directions, decoding semantic information under a global receptive field and linear complexity:
@@
where the expansion and merging operation correspond to VMamba[10], while the S6 operation denotes the selective scan space state sequential model of Mamba[9]. The $\bar{F}$ then goes through layer normalization and an element-wise multiplication with the bypass flow, which has also undergone the same linear embedding and Silu activation. Finally, the VSS block outputs $F_{MSD}$ through a linear layer with residual connection.
@@
II-B Local Supervision Module
Due to the large receptive field of VSS blocks, MSD tends to partially overlook local semantic information while decoding . However, such details in high-resolution images are key factors to further improve the accuracy of semantic segmentation. Therefore, the LSM is proposed to enhance the perception of local details, as depicted in Fig. 1(c).
@@ II-B Local Supervision Module
Considering the relatively smaller scale of land-covers in high-resolution scenarios, we adopt two parallel convolution branches in LSM with the kernel size set as 3 and 1, respectively. Each branch is followed by a batch normalization layer with a ReLU6 activation:
@@
$\tilde{F^{\prime}}=\tilde{F}_{1}^{{}^{\prime}}\oplus\tilde{F}_{3}^{{}^{\prime}}$ (5)
@@
$\tilde{F}=\text{Upsample}\left(\text{Conv}_{1}\left(\text{Dropout}\left(\tilde{F^{\prime}}\right)\right)\right)$ (6)
@@ II-C Loss Functions
To ensure the holistic segmentation performance of UNetMamba, we select two classical loss functions to jointly form our principal loss:
@@
where $L_{dice}$ and $L_{ce}$ represents dice loss and cross-entropy loss, respectively. And $N$ denotes the number of samples while $K$ denotes the number of categories. $\hat{y}_{k}^{\left(n\right)}$ is the $k$-th element in one-hot encoding of the true semantic labels for sample $n$, while $y_{k}^{\left(n\right)}$ is the confidence of sample $n$ belonging to category $k$. Moreover, to achieve the supervisory role of LSM, cross-entropy loss $L_{ce}$ is also selected as the auxiliary loss for local supervision. Therefore, the overall loss function is formulated as follows:
@@
where the weight factor $\alpha$ is set as 0.4 [7].
@@ III-C Performance Comparison
1) Comparison on the LoveDA Dataset: As shown in Tab. I, the proposed UNetMamba achieved the best accuracy performance in quantitative comparison, with a significant improvement of 0.87% in mIoU. It is note-worthy that our UNetMamba achieved SOTA in six out of seven categories. In background and agriculture, UNetMamba led the second place by 1.61% and 2.39%, respectively. This distinct advantage indicated that, benefiting from MSD, our UNetMamba garnered excellent global perception ability, which made it capable of accurately segmenting large land-covers under such a high-resolution condition. The outstanding performance of UNetMamba in building, road and other small land-cover categories proved that the train-only LSM successfully enhanced the perception of local contents through supervision and generalization. The qualitative comparison results on LoveDA are illustrated in Fig. 2, which intuitively shows that the semantic segmentation mask obtained by UNetMamba is more accurate with less omission and clearer boundaries.
2) Comparison on the ISPRS Vaihingen Dataset: The proposed UNetMamba not only achieved SOTA performance on the challenging LoveDA dataset, but also performed noticeably well on the classic Vaihingen dataset. As shown in Tab. II, in terms of accuracy, our UNetMamba achieved improvements of 0.20% in mF1, 0.36% in mIoU, and 0.18% in OA, while maintaining competitive efficiency through a lightweight model (14.76M) with low costs (225.71MB, 100.52G). Fig. 3 illustrates the qualitative comparison results, with the framed areas further demonstrating the competitiveness of UNetMamba in the perception of local semantic details, which is particularly important for the semantic segmentation accuracy in high-resolution remote sensing scenes.
@@ IV Conclusion
In this paper, an efficient semantic segmentation model called UNetMamba is proposed for high-resolution remote sensing images. Considering the multi-scale land-covers and complex information in such images, our UNetMamba features a plug-and-play Mamba-based segmentation decoder MSD for efficient semantic information decoding, and a train-only local supervision module LSM that can efficiently enhance the perception of local semantic details. Extensive experiments conducted on two well-known remote sensing datasets demonstrate that UNetMamba not only outperforms other SOTA models, but also achieves light weight and low costs. In the future, we will continue to explore promising linear mechanisms to further improve the accuracy and efficiency of our UNetMamba.